NERSCPowering Scientific Discovery Since 1974

December 6

Attending:  Eric, Iwona, Mike, Larry, Lisa

Cluster Status/Utilization:  Quite a few free cycles available recently - ALICE/STAR not as busy as usual.

Upcoming Downtimes:  12/13 downtime all day.

Recent Outages/Incidents:  Sunday night the match manager stalled and Jay fixed it.  pdsfdtn1 required a reboot, and there have been some slowness issues related to /common.

Procurements/New Hardware:  All new hardware is in production including pdsfdtn2.  Next procurement will probably be in early 2012.

Shares:  Shares have been calculated and sent to NSD and Physics Division for final adjustments.

SL$ retirement:  Still tentatively scheduled for 3/1/12.

Project Accounts:  In place and working - working on some modifications related to account managment in NIM.

Grid Issues:  Bestman testing on dtn2 is in progress but support is limited.  OSG ony supports gateway bestman now.  This will work for ATLAS but STAR is no longer fully supported.  Iwona will meet with Arie/Alex to discuss future plans.

Jobs in Eqw:  Will adopt new policy of deleting such jobs after being in Eqw state for 5 days.

SL6:  Larry says that we are still on schedule for January rollout.